# Interpretable Early Warnings using Machine Learning in an Online Game-experiment

### v1

@@ Results
Our trained XGBoost model shows predictive power for the time-to-transition up to hours ahead of the transition. For lower values, predictions are closer to the perfect-prediction line (Fig. 4a): at , the mode prediction is ; at , the mode is , and at , it is . As grows, a peak in the distribution of emerges around , corresponding to a bulk of instances the algorithm cannot classify (Fig. 4b). While perfect predictions would sharply peak at for the “no transition” curve, the mode prediction is actually . Nonetheless, performance is determined by the ordering and separation of the predicted values, as predictions are scaled by calibration. The monotonic increase of prediction modes with increasing up to the maximum value is remarkable, as it shows that the algorithm succeeds in discriminating times from transition from times away.
@@
For some compositions, our predictions are close to the perfect-prediction line as a composition approaches transition (Fig. 5a). For compositions where the algorithm performs poorly and oscillates in time (Fig. 5b), there may be continually high activity. When there are no detectable warning signals, should remain relatively constant (Fig. 5c).
To avoid many subsequent warning alarms, which could be inconvenient for a real-world transition, we implement a system with a warning cooldown (see SI section LABEL:secSI:cooldownWarn). Here, the system must “cool down” before issuing another warning, as in Ref. 50. Requirements for such warnings are stricter, which inevitably leads to a lower performance for our cooldown warning system (see the ROC curve in Fig. LABEL:figSI:continuous-warn).
@@
Uncovering mechanisms starts with examining the 172 SHAP trends as a function of feature value (see Figs. LABEL:figSI:SHAPall and LABEL:figSI:SHAPnotime). Our SHAP plots represent time progressing toward the right, so that the right side of the plot (low SHAP) corresponds to feature behaviors that are typical close to transitions. A low SHAP therefore designates instances and features values that the algorithm associates with closeness to transition; a high SHAP (left side of the plot) shifts predictions away from transition. By grouping features that describe the same trend and aspect of dynamics, we build evidence for distinct behaviors which can be considered potential warning signals. To find these behaviors, we first classify the variables for their role in the composition dynamics: strength of the attack or defense changes, image activity, user activity and engagement, image complexity, innovation, and coordination versus noise (see SI section LABEL:secSI:shap for details on the classification). We uncover a set of 12 behaviors of compositions close to transition, shown in Figs. 6 and LABEL:figSI:SHAPinterp.
@@
Trends for past time features provide additional signals. A high past image activity suggests a large presence of past attacks (Fig. LABEL:figSI:SHAPinterpb), which we refer to as a turbulent past. Past high user engagement (Fig. LABEL:figSI:SHAPinterpc) could also indicate substantial past attacks that the defense reacted against. Though a turbulent or highly engaged past signals an approaching transition, low past activity can indicate a neglected composition, which could attract opportunistic attacks (Fig. LABEL:figSI:SHAPinterpd). On the contrary, a very low past image activity may indicate a general lack of interest (Fig. LABEL:figSI:SHAPinterpe), which makes the composition safer.
Attacked pixels left in their new color can indicate evolution of the image supported by the community, which we refer to as innovation. A lack of innovation is indicative of transitions (Fig. 6d), which may suggest that a more innovative community is safer from transition. Coordinated activity rather than noisy activity, whether from the defending community or not, also indicates stronger compositions (Fig. 6e). Unsurprisingly, low defense activity in the past is typical close to transition (Fig. LABEL:figSI:SHAPinterpf), which could mean that the defending community is unable or unmotivated to mobilize. Finally, images of higher complexity are also typically safer (Fig. 6f).
@@ Discussion
This work is, to our knowledge, the first machine-learning-based early warning indicator for transitions in a large-scale human social system. Some machine learning prediction systems using large empirical datasets have outperformed ours in fields such as medicine (49, 50). However, social systems present unique challenges due to the high dimensionality of human behavior and the lack of comprehensive datasets. Decision trees using the r/place dataset tackle both these challenges while providing data-driven insights on the complexity of social dynamics that analytical models cannot offer.
@@
We extract human-readable warning signals from the SHAP analysis of predictions. Many dynamical behaviors, rather than single, generic early warning signals, are found to indicate incoming transitions. These behaviors might apply to transitions in other socio-ecological systems and inform their managers. Some of our findings confirm intuitions from other approaches; for example, innovation has been discussed as contributing to adaptation and sustainability transitions (63), but also as being insufficient preceding the collapse of ancient societies (64, 65). Closer to the theoretical origins of critical slowing down (24, 16), the higher risk of transition for simpler images also evokes the reduction of dimensionality due to power law dynamics close to second-order critical transitions.
@@
These insights capture the diversity of warning signs in a real-world system because we use a large and high-resolution empirical dataset. Forecasting abilities using deep learning have been shown in some ecology (44, 43) and epidemiology (46) studies with training data simulated from simple models inspired by disciplinary tradition. Unlike our algorithm using system-specific time series with features encoding their memory, these methods cannot reveal pre-transition phenomena that are not present in the simulated models. Using empirical data also improves the applicability to related systems, which is illustrated by the striking validity of our 2022 warning system when applied in 2023. Though the r/place experiments of 2022 and 2023 had similar rules, contextual differences altered the 2023 dynamics: for example, differences in expansions of the canvas size and available colors, total duration, cooldown times, moderator and bot activity, level of preparedness of communities, player motivations, and popularity of the game.
@@
While different compositions may show varied warning behaviors, some compositions may show no warnings at all. This cross-subsystem variation is illustrated by the diverse pre-transition behaviors of Figs. 6 and LABEL:figSI:SHAPinterp. It is also supported by recent literature (19), which has proposed a number of indicators of varying generalizability (31). A key advantage of a machine learning warning system is its potential to uncover any warning signals present in real-world data, even when they were not hypothesized in a priori models. Some signals may seem contradictory when seen in the same or similar subsystems, as in the co-occurrence of large critical slowing down and mild critical speeding up in our results. Slowing down typically occurs very close to transition, while speeding up is an earlier, subtle signal, and compositions could show none, one, or both. While slowing down is well-established, speeding up is less conventional but has been discussed (28, 29, 30). Interestingly, a simple logistic growth model (66) displays a direct link between the increase of the reaction rate and the departure from a stable equilibrium. However, these and other analytic indicators all rely on separate, often incompatible models.
@@
The lack of a common framework to describe early warning indicators complicates comparisons between our r/place results and simple indicators from the literature. Different indicators apply to different types of transitions, and identifying the types is not trivial. For example, critical slowing down assumes small perturbations around an equilibrium that slowly moves towards a bifurcation (sometimes called B-tipping), while some transitions are instead triggered by large perturbations that push the state outside the basin of attraction of the initial equilibrium (S- for state-tipping) (67). For sufficiently large and fast perturbations, warning signals may (68) or may not (69, 26) be present. Interestingly, the theory of critical slowing down applies only to second-order, continuous phase transitions and to spinodal instabilities (24), but is not expected in abrupt, first-order transitions that are usually more alarming in real-world systems. Depending on how the transition is modeled, tipping can also be conceived as occurring when the state fails to track a fast-changing equilibrium (rate-induced, or R-tipping) (70); critical slowing down might exist but be delayed in these cases (27). In addition, high-dimensional systems like ours may be more prone to transients, meaning long quasi-stable periods that eventually transition, potentially with no warning signals (37). The transitions in our compositions might be of any type, or even mixed types; for example, coordinated attacks can be seen as a sudden perturbation (S-tipping) or as a fast change of an attack parameter (R-tipping). Our machine learning system does not rely on explicit modeling and can therefore account for this variety of transition dynamics.
@@
More generally, we navigate multiple issues that do not have clear solutions. The first challenge is identifying the equilibrium from which the system departs, including the timescale over which the equilibrium should be evaluated, or even knowing whether an equilibrium exists222When compositions are out of equilibrium, the reference image is only an average of out-of-equilibrium states.. Second, noise, which consists of small perturbations around equilibrium, is difficult to distinguish from coordinated attacks, which actively shift the equilibrium. While our results might in principle depend on these choices, the multiple options we compute often result in highly correlated variables, showing that these options reflect similar dynamics.
@@
Approaching the rate-induced transition, the trend of the variance depends on what pseudo-mean the deviations are computed from, as the mean and equilibrium are ill-defined in our r/place images. The toy model exemplifies repercussions of applying definitions from transition models to variables in a high-dimensional system. The pseudo-mean can be taken as the equilibrium that the state fails to track; but in r/place compositions, this theoretical equilibrium is unknown. Alternatively, the variance can be computed using a short-timescale sample mean—which is only approximate in our compositions, but we use it as the main variance variable. Lastly, the pseudo-mean of the fraction of differing pixels can simply be zero, which corresponds to using the reference image of our compositions. These three variance definitions can lead to different interpretations in the toy model, which underscores the consequences of model design.
@@
Though the r/place data offers unprecedented insights into large-scale human behavior, there are certain limitations. First, we rely on the user-annotated Atlas (59) to classify the pixel change data into compositions. Omissions or inaccuracies from contributors can impact the image borders of compositions, as well as their birth and death times. We rectify some of these mistakes with minimal corrections to composition borders (see SI section LABEL:secSI:compodata); these corrections only impact a small fraction of pixels and do not significantly modify the time series. The collective work of the Atlas contains crucial cultural information that automatic methods such as clustering (14) or edge detection (71) cannot compensate for. Another issue is pixel changes that violated the 5-minute cooldown time. Moderators were not subject to the cooldown rule, and some users had multiple accounts or programmed bots to change pixels automatically (see SI section LABEL:secSI:rplacedata). These pixel changes are hard to eliminate systematically. However, they contribute to the canvas dynamics observed by users and are therefore part of the experiment; they also constitute less than 1% of changes in 2022.
Learning from empirical data also requires caution in the study design; our algorithm would achieve better predictions if we ignored the practicalities of applying a warning system in a real-world scenario. The algorithm can optimize for artifacts that would not be present in a dataset collected live, rather than for genuine features of pre-transition periods. To approximate a live warning system, we: use only information preceding a time instance; use only instances preceded by a stable period; separate whole compositions between train and test samples; and reject compositions that are not yet monitored (see methods section and SI sections LABEL:secSI:transitions and LABEL:secSI:filter). We invite researchers in this field to be wary of showing high-performance predictions at the expense of real-world applicability.
@@
Future work could test these insights in other socio-ecological systems where communities emerge and compete to achieve their goals. In addition, our predictive and interpretable approach is applicable not only to community-driven online platforms, but also to other systems that provide large and dense datasets. Another direction is to train a deep learning algorithm on individual pixel features rather than pixel-averaged features. This approach would bypass the difficulty of defining variables analogous to those in the early warning signals literature and may perform better, though these predictions would be significantly more difficult to interpret. By refining our warning systems and broadening their applicability, our interdisciplinary community can improve detection and understanding of transitions in complex systems.
@@ Conclusion
All figure data is available on the Princeton Data Commons (72). The code reproducing the results of this paper is available on Github at https://github.com/AnnieStephenson/r-place-emergence/tree/EWSpaper.

### v2

@@ Results
Our trained XGBoost model shows predictive power for the time-to-transition up to hours ahead of the transition. For lower values, predictions are closer to the perfect-prediction line (Fig. 4a): at , the mode prediction is ; at , the mode is , and at , it is . As grows, a peak in the distribution of emerges around , corresponding to a bulk of instances the algorithm cannot classify (Fig. 4b). This peak is already evident for . While perfect predictions would sharply peak at for the “no transition” curve, the mode prediction is actually . Nonetheless, performance is determined by the ordering and separation of the predicted values, as predictions are scaled by calibration. The monotonic increase of prediction modes with increasing up to the maximum value is remarkable, as it shows that the algorithm succeeds in discriminating times from transition from times away.
@@
For some compositions, our predictions provide useful warnings where predictions are reasonably accurate near transition (Fig. 5a). For compositions where the algorithm performs poorly and oscillates in time (Fig. 5b), there may be continually high activity. When there are no detectable warning signals, should remain relatively constant (Fig. 5c).
To avoid many subsequent warning alarms, which could be inconvenient for a real-world transition, we implement a system with a warning cooldown (see SI Section LABEL:secSI:cooldownWarn). Here, the system must “cool down” before issuing another warning, as in Ref. hylandEarlyPredictionCirculatory2020. Requirements for such warnings are stricter, which inevitably leads to a lower performance for our cooldown warning system (see the ROC curve in Fig. LABEL:figSI:continuous-warn).
@@
Uncovering mechanisms starts with examining the 175 SHAP trends as a function of feature value (see Figs. LABEL:figSI:SHAPall and LABEL:figSI:SHAPnotime). Our SHAP plots represent time progressing toward the right, so that the right side of the plot (low SHAP) corresponds to feature behaviors that are typical close to transitions. A low SHAP therefore designates instances and features values that the algorithm associates with closeness to transition; a high SHAP (left side of the plot) shifts predictions away from transition. By grouping features that describe the same trend and aspect of dynamics, we build evidence for distinct behaviors which can be considered potential warning signals. To find these behaviors, we first classify the variables for their role in the composition dynamics: strength of the attack or defense changes, image activity, user activity and engagement, image complexity, innovation, and coordination versus noise (see SI Section LABEL:secSI:shap for details on the classification). We uncover a set of 12 behaviors of compositions close to transition, shown in Figs. 6 and LABEL:figSI:SHAPinterp.
@@
Trends for past time features provide additional signals. A high past image activity suggests a large presence of past attacks (Fig. LABEL:figSI:SHAPinterpb), which we refer to as a turbulent past. Past high user engagement (Fig. LABEL:figSI:SHAPinterpc) could also indicate substantial past attacks that the defense reacted against. Though a turbulent or highly engaged past signals an approaching transition, low past activity can indicate a neglected composition, which could attract opportunistic attacks (Fig. LABEL:figSI:SHAPinterpd). In contrast, a very low past image activity may indicate a general lack of interest (Fig. LABEL:figSI:SHAPinterpe), which makes the composition safer.
Attacked pixels left in their new color can indicate evolution of the image supported by the community, which we refer to as innovation. A lack of innovation is indicative of transitions (Fig. 6d), which may suggest that a more innovative community is safer from transition. Coordinated activity rather than noisy activity, whether from the defending community or not, also indicates stronger compositions (Fig. 6e). Unsurprisingly, low defense activity in the past is a transition signal (Fig. LABEL:figSI:SHAPinterpf), which could mean that the defending community is unable or unmotivated to mobilize. Finally, images of higher complexity are also typically safer (Fig. 6f).
@@ Discussion
This work is, to our knowledge, the first machine-learning-based early warning indicator for transitions in a large-scale human social system. Some machine learning prediction systems using large empirical datasets have outperformed ours in fields such as medicine (gaoMachineLearningBased2020, hylandEarlyPredictionCirculatory2020). However, social systems present unique challenges because human behavior depends on many interacting factors and lacks comprehensive datasets. Decision trees using the r/place dataset tackle both these challenges while providing data-driven insights on the complexity of social dynamics that analytical models cannot offer.
@@
We extract human-readable warning signals from the SHAP analysis of predictions. Many dynamical behaviors, rather than single, generic early warning signals, are found to indicate incoming transitions. These behaviors might apply to transitions in other socio-ecological systems and inform their managers. Some of our findings confirm intuitions from other approaches; for example, innovation has been discussed as contributing to adaptation and sustainability transitions (makitieDigitalInnovationsContribution2023), but also as being insufficient preceding the collapse of ancient societies (tainterCollapseComplexSocieties1988, mascarenoCriticalTransitionsEcosystems2022). Closer to the theoretical origins of critical slowing down (levinPhaseTransitionsTheory2023, wisselUniversalLawCharacteristic1984), the higher risk of transition for simpler images also evokes the reduction of dimensionality due to power law dynamics close to second-order critical transitions.
@@
These insights capture the diversity of warning signs in a real-world system because we use a large and high-resolution empirical dataset. Forecasting abilities using deep learning have been shown in some ecology (debMachineLearningMethods2022, buryDeepLearningEarly2021) and epidemiology (miryDeepLearningDisease2025) studies with training data simulated from simple models inspired by disciplinary tradition. Unlike our algorithm using system-specific time series with features encoding their memory, these methods cannot reveal pre-transition phenomena that are not present in the simulated models. Using empirical data also improves the applicability to related systems, which is illustrated by the striking validity of our 2022 warning system when applied in 2023. Though the r/place experiments of 2022 and 2023 had similar rules, contextual differences altered the 2023 dynamics: for example, differences in expansions of the canvas size and available colors, total duration, cooldown times, moderator and bot activity, level of preparedness of communities, player motivations, and popularity of the game.
@@
While different compositions may show varied warning behaviors, some compositions may show no warnings at all. In fact, unpredictable transitions make up a significant part of our data, as shown by the band of poor predictions shown in Fig. 4a and the broadly distributed performance per transition in Fig. LABEL:figSI:ROCb. This cross-subsystem variation is illustrated by the diverse pre-transition behaviors of Figs. 6 and LABEL:figSI:SHAPinterp. It is also supported by recent literature (boettigerEarlyWarningSignals2013), which has proposed a number of indicators of varying generalizability (georgeEarlyWarningSignals2023). A key advantage of a machine learning warning system is its potential to uncover any warning signals present in real-world data, even when they were not hypothesized in a priori models. Some signals may seem contradictory when seen in the same or similar subsystems, as in the co-occurrence of large critical slowing down and mild critical speeding up in our results. Slowing down typically occurs very close to transition, while speeding up is an earlier, subtle signal, and compositions could show none, one, or both. While slowing down is well-established, speeding up is less conventional but has been discussed (titusCriticalSpeedingEarly2020, pomeauCriticalSpeedvsCritical2011, gietkaSqueezingCriticalSpeeding2022). Interestingly, a simple logistic growth model (maySimpleMathematicalModels1976) displays a direct link between the increase of the reaction rate and the departure from a stable equilibrium. However, these and other analytic indicators all rely on separate, often incompatible models.
@@
The lack of a common framework to describe early warning indicators complicates comparisons between our r/place results and simple indicators from the literature. Different indicators apply to different types of transitions, and identifying the types is not trivial. For example, critical slowing down assumes small perturbations around an equilibrium that slowly moves towards a bifurcation (sometimes called B-tipping), while some transitions are instead triggered by large perturbations that push the state outside the basin of attraction of the initial equilibrium (S- for state-tipping) (boettigerBifurcationStateTipping2020). For sufficiently large and fast perturbations, warning signals may (drakeEarlyWarningSignals2013) or may not (boettigerEarlyWarningSignals2012, boettigerNoEarlyWarning2013) be present. Interestingly, the theory of critical slowing down applies only to second-order, continuous phase transitions and to spinodal instabilities (levinPhaseTransitionsTheory2023), but is not expected in abrupt, first-order transitions that are usually more alarming in real-world systems. Depending on how the transition is modeled, tipping can also be conceived as occurring when the state fails to track a fast-changing equilibrium (rate-induced, or R-tipping) (ashwinTippingPointsOpen2012); critical slowing down might exist but be delayed in these cases (ritchieEarlywarningIndicatorsRateinduced2016). In addition, systems like ours with stochasticity, multiple timescales, and more dimensions than typical dynamical models may be more prone to transients, meaning long quasi-stable periods that eventually transition, potentially with no warning signals (hastingsTransientPhenomenaEcology2018). The transitions in our compositions might be of any type, or even mixed types; for example, coordinated attacks can be seen as a sudden perturbation (S-tipping) or as a fast change of an attack parameter (R-tipping). Our machine learning system does not rely on explicit modeling and can therefore account for this variety of transition dynamics.
@@
More generally, we navigate multiple issues that do not have clear solutions. The first challenge is identifying the equilibrium from which the system departs, including the timescale over which the equilibrium should be evaluated, or even knowing whether an equilibrium exists222When compositions are out of equilibrium, the reference image is only an average of out-of-equilibrium states.. We varied the timescale of the equilibrium by training on datasets with different sliding window lengths (see Fig. LABEL:figSI:sensitivitya); the sample of instances and their feature values differ, but the resulting changes in performance are marginal. Second, noise, which consists of small perturbations around equilibrium, is difficult to distinguish from coordinated attacks, which actively shift the equilibrium. While our results might in principle depend on these choices, the multiple options we compute often result in highly correlated variables, showing that these options reflect similar dynamics.
@@
Approaching the rate-induced transition, the trend of the variance depends on which pseudo-mean the deviations are computed from, as the mean and equilibrium are ill-defined in our r/place images. The pseudo-mean can be taken as the equilibrium that the state fails to track; but in r/place compositions, this theoretical equilibrium is unknown. Alternatively, the variance can be computed using a short-timescale sample mean—which is only approximate in our compositions, but we use it as the main variance variable. Lastly, the pseudo-mean of the fraction of differing pixels can simply be zero, which corresponds to using the reference image of our compositions. These three variance definitions can lead to different interpretations in the toy model, which underscores the consequences of model design.
@@
Though the r/place data offers unprecedented insights into large-scale human behavior, there are certain limitations. First, we rely on the user-annotated Atlas (haagmansPlaceAtlasInitiative2024) to classify the pixel change data into compositions. Omissions or inaccuracies from contributors can impact the image borders of compositions, as well as their birth and death times. We rectify some of these mistakes with minimal corrections to composition borders (see SI Section LABEL:secSI:compodata); these corrections only impact a small fraction of pixels and do not significantly modify the time series. The collective work of the Atlas contains crucial cultural information that automatic methods such as clustering (wuLargescaleCollectiveDynamics2024) or edge detection (cannyComputationalApproachEdge1986) cannot compensate for. Another issue is pixel changes that violated the 5-minute cooldown time. Moderators were not subject to the cooldown rule, and some users had multiple accounts or programmed bots to change pixels automatically (see SI Section LABEL:secSI:rplacedata). These pixel changes are hard to eliminate systematically. However, they contribute to the canvas dynamics observed by users and are therefore part of the experiment; they also constitute less than 1% of changes in 2022.
Learning from empirical data also requires careful study design, since prediction performance can be artificially inflated if we ignore the constraints of a real-world warning system. We tested how much removing these precautions increases ROC and PR performance (see SI Section LABEL:secSI:sensitivity and Fig. LABEL:figSI:sensitivityc). First, including compositions that emerge from patchworks—areas where multiple compositions meet and would not be tracked as a single unit—raises performance because the algorithm exploits artifacts that are specific to these patchworks and would therefore not exist in a live setting. Second, if we include the time instances that are too unstable to pass the relative threshold for transitions, the ROC AUC increases because these instances are easily categorized as being far from transition, although the PR AUC decreases due to a substantially smaller signal fraction. In addition, in unstable systems, defining a composition to preserve is ambiguous. Third, using a lower relative threshold of also increases performance, as slower transitions are treated as signal and provide earlier warnings. However, our focus is on abrupt transitions, which are both more difficult and more critical to detect in real-world systems. It is ultimately up to the system manager to decide which systems they wish to preserve against what transitions. In summary, to approximate a live warning system, we reject compositions that are not yet monitored, use only information preceding the assessed time instance, use only instances preceded by a stable period, and separate whole compositions between train and test samples (see Materials and Methods and SI Sections LABEL:secSI:transitions and LABEL:secSI:filter). Without such safeguards, prediction results may show high performance at the expense of real-world applicability.
@@
Future work could test these insights in other socio-ecological systems where communities emerge and compete to achieve their goals. In addition, our predictive and interpretable approach is applicable not only to community-driven online platforms, but also to other systems that provide large and dense datasets. Another direction is to train a deep learning algorithm on individual pixel features rather than pixel-averaged features. This approach would bypass the difficulty of defining variables analogous to those in the early warning signals literature and may perform better, though these predictions would be significantly more difficult to interpret. However, adding complexity in the training dataset with additional spatial variables does not improve our performance (see Materials and Methods and SI Section LABEL:secSI:additionalVars). This might be due to inherent unpredictability of some human behaviors, which is illustrated in Fig. LABEL:figSI:ROCb: transitions in some compositions are easily predicted, while the performance of the algorithm is null for other compositions where transitions could be due to unstrategic, and thus unpredictable, attacks. By refining our warning systems and broadening their applicability, our interdisciplinary community can improve detection and understanding of transitions in complex systems.
@@ Conclusion
All figure data is available on Zenodo (stephenson_anna_2025). The code reproducing the results of this paper is available on Github at https://github.com/AnnieStephenson/r-place-emergence/tree/EWSpaper.